OpenStudy (lovelyharmonics):

okay well i need help with my assignment which is to describe when to use the law of cosines, the law of sines, and the law of sines with the ambiguous case. and i have to provide general guidelines for each law so that it can be applied to any triangle that someones given.... @halorazer

OpenStudy (anonymous):

Yeah, I just did that for you on the other post.

OpenStudy (lovelyharmonics):

no, .-. you gave the definition. i know the definition

OpenStudy (anonymous):

No. I gave an explanation of WHEN to use it.

OpenStudy (anonymous):

You must have missed that.

OpenStudy (lovelyharmonics):

naw brotaco, you gave an explination off google

OpenStudy (anonymous):

No, I seriously didn't. If you didn't know, people actually have to use both of those in math classes? Go ahead and copy and paste my answer and see if you get any results. ._.

OpenStudy (lovelyharmonics):

-.- thats not the way i roll. i just need help with a better explanation than that

OpenStudy (anonymous):

oh okay. should I say a shape with three sides instead of a triangle? is that a better explanation for you, woman?

OpenStudy (lovelyharmonics):

.-.

OpenStudy (anonymous):

;)

OpenStudy (campbell_st):

well for mine, the law of cosines is used when you have 1. 3 sides of a triangle and needs to find an angle you use \[\cos(A) = \frac{ b^2 + c^2 - a^2}{2bc}\] or 2. You have 2 sides and the included angle and then need to find a 3rd side you use \[a^2 = b^2 + c^2 - 2\times b \times c \times \cos(A)\] for the law of sines you can be given 1. 2 sides and an angle...(thats not the included angle). Where you need to find an angle. \[\frac{\sin(A)}{a} = \frac{B}{b}\] 2. 2 angles and you need to find a side \[\frac{a}{\sin(A)} = \frac{b}{\sin(B)}\] with the side rule taking the reciprocals of each fraction makes it easy to find the missing information. The ambiguous case occurs when the angle is obtuse... this needs a little checking first.... and remember in triangles the smallest angle is opposite the shortest side the middle angle is opposite the middle length side the largest angle is opposite the longest side... hope it helps

OpenStudy (lovelyharmonics):

yeah that does c: thank you

OpenStudy (campbell_st):

oops... 1st law of sines is \[\frac{\sin(A)}{a} = \frac{\sin(B)}{b}\]